# Streamlining Redundant Layers to Compress Large Language Models

### v3

@@ 3.2 Stability Metric
We propose a novel metric “stability” to evaluate the performance of LLMs after pruning, i.e.,
@@
where the identifier $\mathbbm{1}_{\left[i\in{\rm TP}\cup{\rm TN}\right]}$ is used to indicate whether the $i$-th sample belongs to ${\rm TP}$ and ${\rm TN}$. We use ${\rm std}_{i}$ as the weight of the $i$-th sample. Because the ${\rm std}$ of different samples varies significantly. to mitigate the influence of samples with excessively large standard deviations, we apply the $\exp$ function to moderate the weight differences among samples. Different from accuracy, stability focuses on the model’s confidence in its answers and the consistency between the model before and after pruning on a task, aligning more closely with the goal of model pruning, i.e., ensuring the pruned model remains as similar as possible to the original model.
@@ 4 Experiments
In this section, we first compare our proposed method, LLM-Streamline, with several popular structured pruning methods to demonstrate the effectiveness of our method (4.4). We also highlight the necessity of the lightweight network by ignoring it. After that, we analyze the impact of different sizes and structures of lightweight networks on the model’s performance (4.5) and evaluate the performance under various pruning ratios (4.6). Finally, we compare the proposed layer replacement with the well-known PEFT method, LoRA hu2021lora (4.7), demonstrating that layer replacement provides superior performance and reduced memory overhead.
@@ 4.1 Setup
To validate the effectiveness of our proposed model, we conduct experiments on popular open-source LLMs, including OPT-6.7Bzhang2022opt and Llama2-7B touvron2023llama . Following previous work, we perform experiments with a 25% compression rate and extract data from SlimPajama for layer pruning and layer replacement, where SlimPajama is a dataset containing data from different domains for pre-training LLM. Sheared LLaMa xia2023sheared finds that the performance degradation of pruned models varies across different domains, and proposes determining the distribution of data from different domains based on the degree of performance degradation. Therefore, we randomly sample the data based on the distribution used by Sheared LLaMa xia2023sheared , finally constructing the dataset containing 30,000 pieces of data. We randomly select 500 samples from this dataset and input them into LLMs, generating Fig. 2, and use these 500 data samples for layer pruning. All 30,000 pieces of data are used to train the lightweight network. We use two types of lightweight networks: Feed-Forward Neural Network (FFN) and Transformer Layer. The FFN is randomly initialized, while the Transformer Layer inherits the parameters of the first pruned layer. When training the FFN and Transformer layers, different learning rates are used: $10^{-3}$ for the FFN and $10^{-5}$ for the Transformer layer. Both are trained for $20$ epochs with a batch size of $32$. On an A800 GPU, the training process takes approximately $5$ hours to complete. Further experimental details are available in the Appendix B.1.
@@ 4.2 Benchmark
To comprehensively assess the changes in LLM capabilities before and after pruning, we follow the evaluation method of LaCo yang2024laco . We use 12 natural language understanding and question-answering benchmarks for evaluation: CMNLI xu2020clue , HellaSwag(HeSw) zellers2019hellaswag , PIQA bisk2020piqa ,CHID zheng2019chid , WSC levesque2012winograd ,CommonSenseQA(CoQA) talmor2018commonsenseqa , BoolQ clark2019boolq ,MMLU hendrycks2020mmlu , CMMLU li2023cmmlu ,Race-High/Middle lai2017race , C3 sun2020C3 . The tasks in these benchmarks are formalized as classification tasks, so we refer to these benchmarks as classification benchmarks. For these benchmarks, we use both accuracy and stability as metrics for evaluating the models. Additionally, we include 3 benchmarks: XSum narayan2018xsum , GSM8K cobbe2021gsm8k and StrategyQA geva2021did , to demonstrate the LLM’s performance on generation tasks after pruning. We refer to these tasks as generation benchmarks. Following the evaluation framework of OpenCompasscontributors2023opencompass , we use accuracy to evaluate StrategyQA and GSM8K, and use ROUGE1 to evaluate Xsum.
@@ 4.3 Baseline
We compare several structured pruning methods that prune the attention head, the filters of the FFN layer, and the hidden dimension. We also compare existing layer-pruning methods.
LLM-Prunerma2023llmpruner prunes attention heads, FFN layer filters, and hidden dimensions by using gradients and activations to estimate the importance of these modules.
SliceGPTashkboos2024slicegpt prunes hidden dimensions. It inserts dimensionality reduction matrices into the model and employs Principal Component Analysis (PCA) to initialize and eliminate unimportant components within these matrices. The reduced-dimensional matrices are subsequently merged with the model’s original weight matrices, thereby reducing the model’s size.
LaCoyang2024laco prunes layers by dividing the layers into groups, each consisting of multiple consecutive layers, and compress them separately, whereas our method simply compresses a piece of consecutive layers. LaCo merges consecutive layers by averaging their parameters whereas we train an additional lightweight network to replace these layers.
LLM Method Ratio Benchmarks Ave. Per.
@@
Table 2: Accuracy of pruning methods on classification benchmarks. “*” indicates that we refer to the results in the original paper. ShortGPT and UIDL men2024shortgpt ; gromov2024unreasonable can be considered as a special case of our approach, i.e., Ours (None). Preformance (Pre.) represents the percentage of the original model’s performance retained by the pruning method.
LLM Method Ratio Benchmarks Ave.
@@
LLM Method Ratio Benchmarks Ave. Per.

### v4

@@ 3.2 Stability Metric
We propose a novel metric stability to evaluate the performance of LLMs after pruning, i.e.,
@@
where the identifier $\mathbbm{1}_{\left[i\in{\rm TP}\cup{\rm TN}\right]}$ is used to indicate whether the $i$-th sample belongs to ${\rm TP}$ and ${\rm TN}$. We use ${\rm std}_{i}$ as the weight of the $i$-th sample. Because the ${\rm std}$ of different samples varies significantly, to mitigate the influence of samples with excessively large standard deviations, we apply the $\exp$ function to moderate the weight differences among samples. Different from accuracy, stability focuses on the model’s confidence in its answers and the consistency between the model before and after pruning on a task, aligning more closely with the goal of model pruning, i.e., ensuring the pruned model remains as similar as possible to the original model.
@@ 4 Experiments
In this section, we first compare our proposed method, LLM-Streamline, with several popular pruning methods to demonstrate its effectiveness (4.4). Next, we analyze the impact of different sizes and structures of lightweight networks on model performance (4.5) and evaluate performance under various pruning ratios (4.6). Finally, we compare our layer replacement approach with the well-known PEFT method, LoRA (Hu et al., 2021) (4.7), showing that layer replacement offers superior performance and reduced memory overhead.
@@ 4.1 Setup
To validate the effectiveness of our proposed method, we conduct experiments on popular open-source LLMs, including Llama2-7B and Llama2-13B  (Touvron et al., 2023). Following previous work (Men et al., 2024; Yang et al., 2024), we perform experiments with a 25% compression rate and extract data from SlimPajama for layer pruning and layer replacement, where SlimPajama is a dataset containing data from different domains for pre-training LLM. Sheared LLaMa  (Xia et al., 2023) finds that the performance degradation of pruned models varies across different domains, and proposes determining the distribution of data from different domains based on the degree of performance degradation. Therefore, we randomly sample the data based on the distribution used by Sheared LLaMa  (Xia et al., 2023), finally constructing the dataset containing 30,000 pieces of data. We randomly select 500 samples from this dataset and input them into LLMs, generating Fig. 2, and use these 500 data samples for layer pruning. All 30,000 pieces of data are used to train the lightweight network. We utilize two types of lightweight networks: a Feed-Forward Neural Network (FFN), referred to as Ours (FFN), and a Transformer Layer, referred to as Ours (Layer). The FFN is randomly initialized, while the Transformer Layer inherits the parameters from the first pruned layer. Additionally, we explore a purely pruning approach without incorporating any lightweight network, denoted as Ours (None). Further experimental details are available in the Appendix C.1.
@@ 4.2 Benchmark
To comprehensively assess the changes in LLM capabilities before and after pruning, we use 12 natural language understanding and question-answering benchmarks for evaluation: CMNLI  (Xu et al., 2020), HellaSwag(HeSw)  (Zellers et al., 2019), PIQA  (Bisk et al., 2020),CHID  (Zheng et al., 2019), WSC  (Levesque et al., 2012),CommonSenseQA(CoQA)  (Talmor et al., 2018), BoolQ  (Clark et al., 2019),MMLU  (Hendrycks et al., 2020), CMMLU  (Li et al., 2023),Race-High/Middle  (Lai et al., 2017), C3  (Sun et al., 2020). The tasks in these benchmarks are formalized as classification tasks, so we refer to these benchmarks as classification benchmarks. For these benchmarks, we use both accuracy and stability as metrics for evaluating the models. Additionally, we include 3 benchmarks: XSum  (Narayan et al., 2018), GSM8K  (Cobbe et al., 2021) and StrategyQA  (Geva et al., 2021), to demonstrate the LLM’s performance on generation tasks after pruning. We refer to these tasks as generation benchmarks. Following the evaluation framework of OpenCompass (Contributors, 2023), we use accuracy to evaluate StrategyQA and GSM8K, and use ROUGE1 to evaluate Xsum.
@@ 4.3 Baseline
We compare several pruning methods that prune the attention heads, the filters of the FFN layer, and the hidden dimension, as well as the concurrent layer-pruning methods LaCo. In addition, ShortGPT and UIDL (Men et al., 2024; Gromov et al., 2024) can be considered as a variant of our approach, i.e., Ours (None). We also discuss layer pruning methods which use perplexity as the metric, such as SLEB (Song et al., 2024), in Appendix A.
LLM-Pruner (Ma et al., 2023) prunes attention heads, FFN layer filters, and hidden dimensions by using gradients and activations to estimate the importance of these modules.
SliceGPT (Ashkboos et al., 2024) prunes hidden dimensions. It inserts dimensionality reduction matrices into the model and employs Principal Component Analysis (PCA) to initialize and compress the matrices, and then merge them with the original weight matrix to reduce the model’s size.
LaCo (Yang et al., 2024) prunes layers by dividing the layers into groups, each consisting of multiple consecutive layers, and compresses them separately, whereas our method simply compresses a piece of consecutive layers. LaCo merges consecutive layers by averaging their parameters whereas we train an additional lightweight network to replace these layers.
LLM Method Ratio Benchmarks Average RP
@@
Llama2-13B Dense 0.00% 47.5 33.0 47.2 71.5 51.0 66.8 74.8 79.8 60.0 58.1 55.8 38.7 57.0 100.0
LLMPruner 24.4% 29.5 33.0 29.5 58.0 47.1 43.7 54.7 72.7 21.9 22.5 25.2 24.9 38.6 67.7
SliceGPT 23.6% 38.6 30.5 18.3 37.8 42.3 38.3 45.6 61.9 24.0 25.0 30.6 25.6 34.9 61.2
LaCo* 24.6% 44.9 32.9 40.1 64.0 52.9 52.7 64.4 74.3 56.6 54.5 45.9 32.6 51.3 90.0
Ours (None) 24.6% 47.0 33.0 36.5 62.3 64.4 58.8 66.6 73.5 60.2 58.3 54.8 38.4 54.5 95.6
Ours (FFN) 25.4% 45.8 33.0 37.1 67.4 37.5 64.4 67.9 74.0 58.6 58.2 55.7 38.6 53.2 93.3
Ours (Layer) 24.6% 45.7 33.0 38.0 66.2 36.5 63.8 69.1 75.1 58.0 57.4 55.1 39.2 53.1 93.2
Table 2: Accuracy of pruning methods on classification benchmarks. “*” indicates that we refer to the results in the original paper. Retained performance (RP) represents the percentage of the original model’s performance retained by the pruning method.
LLM Method Ratio Benchmarks Average
@@
Llama2-13B LLMPruner 24.4% 71.6 100 69.2 70.5 65.4 69.5 77.8 86.7 42.3 35.6 48.1 52.3 65.8
SliceGPT 23.6% 62.2 39.5 51.4 27.1 68.3 65.5 64.9 75.6 45.3 43.4 52.7 52.9 54.1
Ours (None) 24.6% 84.2 99.9 71.8 77.4 46.2 82.2 85.7 86.5 83.3 83.6 89.1 83.8 81.1
Ours (FFN) 25.4% 85.7 100 72.5 79.8 59.6 89.2 89.4 89.7 84.8 83.3 93.6 90.7 84.9
Ours (Layer) 24.6% 87.4 100 74.1 81.3 58.6 89.0 90.5 90.5 84.2 83.0 92.5 85.5 84.7
@@
LLM Method Ratio Benchmarks Average RP
@@
Llama2-13B Dense 0.00% 23.7 29.0 58.1 36.9 100.0
LLMPruner 24.4% 17.5 1.9 43.7 21.0 56.9
SliceGPT 23.6% 5.0 1.9 38.3 15.1 40.9
LaCo* 24.6% 14.5 - - - -
Ours (None) 24.6% 17.7 2.35 46.0 22.0 59.6
Ours (FFN) 25.4% 21.4 4.10 59.6 28.4 77.0
Ours (Layer) 24.6% 21.8 4.70 57.3 27.9 75.6
Accuracy 45.1 45.2 45.6 44.4 46.0 45.8 43.8 44.2
Stability 81.2 82.7 81.9 79.2 80.7 83.7 82.6 83.3
[…]
4.4 Main Results
We present accuracy and stability for different methods on the classification benchmarks in Table 2 and Table 3, respectively, and Table 4 for the generative benchmarks. The results demonstrate that our proposed LLM-Streamline consistently outperforms the baseline methods, effectively retaining the majority of the capabilities of LLMs across various tasks. Specifically, in classification tasks, LLM-Streamline surpasses LLM-Pruner by 7% in accuracy and 12% in stability on Llama2-7B, and by 16% in accuracy and 19% in stability on Llama2-13B. LLM-Streamline also surpasses LaCo by 5% in accuracy on Llama2-7B. For generation tasks, LLM-Streamline retains nearly 77% of Llama2-7B and Llama2-13B’s capabilities, significantly outperforming other pruning methods. We find that almost all of the pruning methods fail on the GSM8K dataset, which reflects the difficulty of math word problems. However, we find that sufficient training can gradually restore the model’s performance on math tasks, and the specific experimental results are shown in Table 21 of Appendix D.5.
Additionally, comparing the average stability (Average) in Table 3 with the retrained performance (RP) in Table 2 across different pruning methods reveals that stability is often much lower than accuracy. We also observe that accuracy on Race-M and Race-H even increases after model pruning. Furthermore, we find that without using any lightweight network, Llama2-13B achieves the highest accuracy on classification benchmarks, but its stability on classification benchmarks and performance on generation benchmarks are lower. These results indicate that pruned models tend to make correct guesses on some classification questions that they are uncertain, highlighting the limitations of accuracy as a sole measure of pruning method performance. We also conduct experiments on OPT-1.3B, OPT-2.7B, OPT-6.7B, Baichuan-7B, Baichuan-13B, Baichuan2-7B and Baichuan2-13B. Details can be found in Appendix D.1, Appendix D.2, Appendix D.3 and Appendix D.4.
4.5 Impact of Different Lightweight Networks
While FFN achieves the best result, Transformer layer still has performance potential. We perform experiments with Llama2-7B using various lightweight network architectures, including Feed-Forward Neural Networks (FFN), SwiGLU-based Feed-Forward Neural Networks (SwiGLU), and Transformer layers. We also explore various initialization methods for the Transformer Layer, including random initialization (Layer-Random), inheritance of the first pruned layer (Layer-First), inheritance of the last pruned layer (Layer-Last), and averaging the pruned layers (Layer-Avg). The average accuracy and stability metrics across all the classification benchmarks are presented in Table 5, with detailed results on each benchmark available in Appendix D.6. The results show that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheriting the pruned first layer yields the best results. In contrast, the performance of Layer-Avg, inspired by LaCo, shows that averaging weights does not achieve the same effectiveness as the pruned first layer.
In addition, we plot the validation loss curves during the training process for different lightweight networks, as shown in Fig. 3. We can observe that FFN and SwiGLU have already converged by the 10th epoch, whereas the loss of Transformer Layer is still decreasing. This indicates that the Transformer layer still has potential, and further training could yield better results, but this would require more computing resources.
4.6 Impact of Different Pruning Ratios
The performance of the pruned model is linearly correlated with the number of parameters. To verify the model’s performance at different pruning ratios, we evaluate our method not only at the approximately 25% pruning ratio but also at ratios of around 2%, 11%, and 20% on Llama2-7B. The average stability and accuracy metrics across all the classification benchmarks are shown in Fig. 4, with details on each benchmark presented in Appendix D.7. By comparing the performance of the original Llama2-7B, TinyLlama-1.1B, OpenLlama-3B-v2, and Llama2-7B pruned at various ratios, we observe a linear correlation between the performance of both the pruned models and the pre-trained original models relative to the number of parameters. This suggests that the performance of models pruned using our method is comparable to that of pre-trained models with the same number of parameters.
4.7 Comparison of Layer Replacement and LoRA
Layer Replacement outperforms LoRA in both performance and GPU memory consumption. We compare the performance of layer replacement with LoRA. Since layer replacement is trained based on hidden states with a different training objective than LoRA, we additionally train one epoch using the language model loss for layer replacement when comparing it with LoRA. The training details can be found in the Appendix C.2. For layer replacement, we freeze the original model’s weights and train only the lightweight network. In the case of LoRA, we set the rank to 128 to align the number of parameters trained with those of the lightweight networks. We randomly extract 30,000 samples from SlimPajama-6B for layer replacement training and also test with the entire dataset to evaluate the limited impact of extensive data on performance (details in Appendix D.5). For LoRA, we use 300,000 samples from SlimPajama-6B. Table 6 presents the average accuracy and stability across all classification benchmarks, with detailed results available in Appendix D.8. The findings indicate that layer replacement surpasses LoRA in both accuracy and stability, while also requiring significantly less GPU memory and training data.
Figure 4: (a) Stability of the pruned Llama2-7B at pruning ratios of approximately 2%, 11%, 20%, and 25%. Stability is calculated only for the compressed model compared to the original, so it cannot be computed for TinyLlama-1.1B and OpenLlama-3B-v2. (b) Accuracy of the pruned Llama2-7B at different pruning ratios, compared to the original Llama2-7B, OpenLlama-3B-v2, and TinyLlama-1.1B. Metrics are averaged across classification benchmarks.
Layer-First Layer-Last Layer-Avg FFN FFN SwiGLU SwiGLU LoRA
Accuracy 46.7 46.8 46.2 45.8 46.3 44.4 45.5 44.5
Stability 85.7 85.6 83.9 83.4 85.2 84.7 84.7 82.1
GPU Memory (G) 27.8 27.8 27.8 25.6 27.0 25.3 26.4 56.4
[…]
Method Metric Need Training Training Data Data Size Training Module Trainig Method
SLEB Perplexity No None None None None
ShortGPT Cosine Similarity No None None None None
UIDL Cosine Similarity Yes C4 164M LoRA-Adapter QLoRA
LaCO Cosine Similarity Yes Unpublished 1B Full Parameters Fine-tuning
Shortened Llama Taylor Perplexity Yes SlimPajama Alpaca 627B 50k Full Parameters LoRA-Adapter Fine-tuning LoRA
LLM-Streamline Cosine Similarity Yes SlimPajama 30k Lightweight Network Training Lightweight Network
Table 7: Comparison of concurrent layer pruning methods, with the metric indicating the importance of layers. Shortened Llama consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
5 Related Work
Previous pruning methods for LLMs primarily focus on pruning dense matrices (Ashkboos et al., 2024), attention heads (Michel et al., 2019; Voita et al., 2019), filters (McCarley et al., 2019; Prasanna et al., 2020), or hidden dimension (Xia et al., 2023; van der Ouderaa et al., 2023). These approaches often lead to structural irregularities, making pruned models less flexible for deployment. In contrast, layer pruning, which only alters the model’s depth, is easier to deploy. Concurrent works in layer pruning alongside LLM-Streamline include LaCo (Yang et al., 2024), ShortGPT (Men et al., 2024), UIDL (Gromov et al., 2024), SLEB (Song et al., 2024), and Shortened Llama (Kim et al., 2024).
LaCo (Yang et al., 2024) divides layers into groups of consecutive layers and compresses them by replacing the consecutive layers with averaged parameter weights. ShortGPT (Men et al., 2024) uses a BI score, equivalent to cosine similarity, to assess layer importance and remove less important layers. Similarly, UIDL (Gromov et al., 2024) uses angular distance, also equivalent to cosine similarity, to determine and remove less important layers, and employs QLoRA to enhance performance. SLEB (Song et al., 2024) calculates layer importance using perplexity and discards those deemed insignificant. Shortened Llama (Kim et al., 2024) explores various layer selection metrics and examines the effectiveness of using continual pre-training and LoRA after pruning. The differences between these layer pruning methods and LLM-Streamline are summarized in Table 7.
Unlike traditional layer pruning methods, LLM-Streamline fundamentally differs by retraining a lightweight model to replace the pruned layers, rather than removing them directly with or without training the pruned model. LLM-Streamline reduces both computation time and resource consumption compared to layer pruning methods (Shortened Llama, LaCo, UIDL) that necessitate retraining. Additionally, LLM-Streamline better preserves the performance of the original LLM compared to concurrent layer pruning methods.
6 Conclusion
In this paper, we propose LLM-Streamline, a layer pruning-and-replacement algorithm for LLMs. We also identify shortcomings in the existing accuracy metric and propose a new metric called stability for evaluating model compression. Extensive experiments show that this layer replacement method using a lightweight network outperforms previous state-of-the-art pruning methods and demonstrates superior effectiveness and efficiency compared to concurrent layer pruning methods.
Appendix A Comparison of Cosine Similarity and Perplexity
To demonstrate the sensitivity of perplexity, referencing the SLEB (Song et al., 2024), we prune the Llama2-7B model with different pre-training datasets, including SlimPajama, C4 and wikitext. The experimental results are presented in Table 8. When pruning with cosine similarity, the layers pruned are consistent across different datasets, whereas when pruning with perplexity, the layers vary, indicating the sensitivity of perplexity. In addition, we evaluate the model after pruning with the SlimPajama dataset, and the experimental results are shown in Table 9. This indicates that the model pruned with perplexity shows lower perplexity on the dataset used for pruning, but performs worse on downstream tasks.
LLM Dataset Pruned Layers
Perplexity Cosine Similarity
Llama2-7B SlimPajama 9,10,11,12,21,23,25,27 22,23,24,25,26,27,28,29
wikitext 9,10,11,12,21,23,24,27 22,23,24,25,26,27,28,29
C4 8,9,11,12,22,23,24,25 22,23,24,25,26,27,28,29
Table 8: Pruned layers using perplexity and cosine similarity for pruning.
LLM Metric Perplexity* Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU Xsum GSM8k StrategyQA
Llama2-7B Dense 6.23 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 19.4 16.5 60.2 45.7 100.0
Cosine Similarity 19.7 40.2 34.4 21.5 67.3 40.4 51.7 59.7 69.0 35.5 34.7 44.6 28.9 14.8 1.97 41.8 39.1 85.6
Perplexity 12.1 37.6 33.0 34.2 61.7 36.5 47.3 56.5 71.4 22.1 21.6 25.9 24.8 17.1 1.74 33.2 35.0 76.6
Table 9: Detailed results of accuracy of using perplexity and cosine similarity for pruning. “Perplexity*” refers to the Perplexity of the pruned model on SlimPajama. Using perplexity as the metric can be considered as SLEB, while using cosine similarity as the metric can be considered as a variant of our approach, i.e., Ours (None)(details in Section 4.1).
Appendix B Data Distribution
We extract the training data from different domains based on the data distribution strategy proposed in Sheared-LLaMa  (Xia et al., 2023). The detailed data distribution is shown in Table 10.
CC GitHub Book StackExchange Wiki ArXiv C4
SlimPajama-6B 54.1% 4.2% 3.7% 2.8% 3.1% 3.4% 28.7%
Ours 36.1% 0.8% 9.1% 1.0% 3.1% 0.7% 49.2%
Appendix C Training Implementation Details
C.1 Lightweight Network Training Details
C.2 Post Training Details
For layer replacement, in order to have a fairer comparison with LoRA, we conduct one epoch of post-training with a learning rate of 5e-5, a weight decay of 1e-3, and a batch size of 32. This process takes less than an hour on a single A800 GPU. For LoRA, the model is trained one epochs with a learning rate of 1e-4, a weight decay of 1e-3, and a batch size of 32. Since the amount of training data used is ten times that of layer replacement, it take approximately 10 hours to complete the training on a single A800 GPU.
Appendix D Detailed Experimental Results
D.1 Experimental Results of OPT-6.7B
We also conduct experiments on OPT-6.7B. The experimental results are shown in Table 11, Table 12 and Table 13. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Table 11: Accuracy of different pruning methods on classification benchmarks by pruning OPT-6.7B.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Table 12: Stability of different pruning methods on classification benchmarks by pruning OPT-6.7B.
LLM Method Ratio Benchmarks Average RP
Xsum GSM8K StrategyQA